Prove or disprove that the two propositions in each pair are equivalent.

 (p "\bigoplus" (q "\to" r)), ((p "\bigoplus" q) "\to" (p "\bigoplus" r))


1
Expert's answer
2022-03-27T16:07:42-0400

"p=q=r=1:\\p\oplus \left( q\rightarrow r \right) =1\oplus 1=0\\\left( p\oplus q \right) \rightarrow \left( p\oplus r \right) =0\rightarrow 0=1\\Not\,\,equivalent"